# Tomography of high-twist proton structure through $ep$ elastic scattering

### v1

@@ Abstract
We present the first high-twist study of the proton form factors $F_{1,2}(Q^{2})$ in $ep$ elastic scattering based on the perturbative QCD $k_{T}$ factorization, $Q^{2}$ being momentum transfer squared. It is motivated by unexpectedly large higher-power contributions from subleading-twist light-cone distribution amplitudes (LCDAs), which are attributed to the enhancement in endpoint regions of parton momentum fractions. We highlight that the endpoint enhancement, tamed by the $k_{T}$ resummation effect, is crucial for accommodating the approximate scaling behavior of the $Q^{4}F_{1}(Q^{2})$ data at intermediate $Q^{2}\sim{\cal O}(10)$ GeV2. The proton LCDAs up to twist 6 are then extracted, and verified by the charge-parity asymmetries observed in the heavy baryon decays $\Lambda_{b}\to p\pi,pK$. Our work provides new insights into the proton three-dimensional structure and manifests the precision requirement for reliable perturbative analyses of baryonic exclusive processes.
@@
We present the first pQCD study of the proton FFs at subleading powers, including the LCDAs up to twist 6 systematically. It will be shown that the endpoint enhancement maximizes the contributions to the FFs at combined twist 8, and the power expansion starts to converge above combined twist 9. The match of the pQCD predictions to the precise FF data from SLAC Arnold:1986nq ; Sill:1992qw ; Rock:1991jy and JLab Christy:2021snt selects the preferred parameters in the proton LCDAs, which confront the outcomes from QCD sum rules (QCDSR) Chernyak:1987nt , lattice QCD (LQCD) RQCD:2019hps and data-driven light-cone sum rules Anikin:2013aka . The determined proton LCDAs, which ought to be universal, are verified by the improved consistency between the predictions and the data for the charge-parity violation in the heavy baryon decays $\Lambda_{b}\to p\pi,pK$ Han:2024kgz ; Han:2025tvc ; LHCb:2025ray . We highlight that the endpoint enhancement, tamed by the $k_{T}$ resummation effect, is crucial for explaining the scaling of $Q^{4}F_{1}(Q^{2})$, which cannot be achieved in the QCDF Huang:2024ugd ; Chen:2024fhj . Our work elaborates the necessity of the proton three-dimensional structure at high twists for understanding the $ep$ elastic scattering data.
@@
in which the plus (minus) component reads $P^{+}=E+Q/2$ ($P^{-}=E-Q/2$) with the proton energy $E=\sqrt{Q^{2}/4+m_{p}^{2}}$, and the parton momentum fractions $x_{i}^{(\prime)}$ and transverse momenta ${\bf k}^{(\prime)}_{iT}$ obey the conservations $\sum_{i}x^{(\prime)}_{i}=1$ and $\sum_{i}{\bf k}^{(\prime)}_{iT}=0$.
@@
The parton transverse momenta span three hierarchical scales: $Q$ (hard), $\sqrt{Q\Lambda_{\rm QCD}}$ (hard-collinear) and $\Lambda_{\rm QCD}$ (soft). The infrared logarithms associated with the soft scale are absorbed into the wave functions $\psi$. The large logarithms associated with the hard-collinear scale are resummed into the well-known Sudakov factors in the impact parameter space, and factored out of $\psi$, $\psi=e^{-S(x_{i},b_{i},Q)}\phi(x_{i},b_{i},\mu_{f})$, $\phi$ being the proton LCDAs. The explicit expression for the exponent $S$ have been given in Refs. Li:1992ce ; Botts:1989kf ; Li:1992nu . The scales $\mu_{r}=\mu_{f}$ for each diagram are set to the maximal virtuality of internal quarks and gluons. With the soft and hard-collinear dynamics being organized appropriately, Eq. (3) is applicable to exclusive processes at intermediate energies; the low $k_{T}$ (high $b$) region would be strongly suppressed by the Sudakov factors, such that internal particle virtualities are governed by finite transverse momenta in the endpoint region. We will ignore the intrinsic $b$ dependence and the factorization scale $\mu_{f}$ in $\phi(x_{i},b_{i},\mu_{f})$ for simplicity, so that the parton transverse distributions are completely described by the Sudakov factors.
@@
respectively, where $C$ stands for the charge conjugation, and $z$ is a light-like vector chosen in the minus direction. At NLO of the conformal spin expansion, five additional dimensionless parameters are related to local operators containing one derivative: $V_{1}^{d}$, $f_{1}^{d}$, $f_{2}^{d}$ ($A_{1}^{u}$, $f_{1}^{u}$) are associated with the derivatives on the $d$-quark ($u$-quark) fields in Eq. (4). The precise information on the above parameters are not yet available. Results from QCDSR Braun:2000kw ; Braun:2006hz and LQCD RQCD:2019hps , both acquired by the Regensburg group, exhibit magnificent discrepancies and uncertainties, even for the leading conformal spin parameters $f_{p}$, $\lambda_{1}$, and $\lambda_{2}$. Alternative methods for determining these parameters, such as data-driven analyses of the FFs, are worthwhile and can yield critical insights into the proton structure.
Table 1: Typical integrands of $F_{1}(Q^{2})$.
@@
Figure 1 presents the $Q^{2}$ dependencies of the contributions from different twists to $Q^{4}F_{1}(Q^{2})$ with the QCDSR inputs of the nonperturbative parameters Braun:2006hz , where only those from even twists are shown for simplicity. It turns out that the pieces from twist $n=8$ dominate, and the power expansion converges above $n=9$. Comparing Fig. 1 with the corresponding data in Fig. 2(a), we see that the predicted $Q^{2}$ dependence deviates dramatically from the observed scaling within $5<Q^{2}<30$ GeV2, though the magnitudes of $Q^{4}F_{1}(Q^{2})$ are of the same order. We stress that the roughly constant $Q^{4}F_{1}(Q^{2})$ around intermediate $Q^{2}$ cannot be understood in collinear factorization, since the dominant $1/\tilde{Q}^{8}$ contribution leads to the descent of $Q^{4}F_{1}(Q^{2})$ with $Q^{2}$ definitely. Considering the additional hard-collinear scale $k_{T}$, which may modify the power-law behaviors associated with different twists, we examine this issue in the pQCD approach.
@@
If two parameters are always present as a product in the formula, the product is chosen as an element of $\left[a_{i}\right]$. The hard matrix ${\cal H}_{ij}$ depends only on $Q^{2}$, which has absorbed the scale evolution of the parameters. The pQCD predictions for the reduced $ep$ cross section at selected $Q_{l}^{2}$ points $\sigma_{l}^{\rm pQCD}\equiv\sigma^{\rm pQCD}(Q_{l}^{2})$ are obtained by inserting Eq. (5) into Ref. Christy:2021snt
@@
QCDSR $5.0\pm 0.5$ $(2.7\pm 0.9)$ $5.4\pm 1.9$ $0.23\pm 0.03$ $0.38\pm 0.15$ $0.4\pm 0.05$ $22\pm 0.05 0.0$$07\pm 0.05$
LQCD $3.67\pm 0.06$ $(4.02\pm 0.38)$ $8.37\pm 0.43$ $288\pm 0.007 0.0$$096\pm 0.010$
this work $3.67\pm 0.06$ $(3.98\pm 0.37)$ $8.32\pm 0.43$ $1.15\pm 0.44$ $1.27\pm 0.72$ $0.85\pm 0.12$ $0.59\pm 0.04$ $(0.23\pm 0.03)$
The parameters from the fit with $\chi^{2}=14$ are compared with those from QCDSR Braun:2006hz and LQCD RQCD:2019hps in Table 2. The fit stabilizes the prior set of $f_{p}$, $\lambda_{1}$ and $\lambda_{2}$ derived in LQCD, including their uncertainties, confirming that these values are reasonable. However, significant shifts in $V_{1}^{d}$, $A_{1}^{u}$, $f_{1}^{u}$, $f_{1}^{d}$ and $f_{2}^{d}$ relative to QCDSR and LQCD results arppear. Because our pQCD study is conducted at LO in $\alpha_{s}$, the distinction can be conclusive after higher-order corrections are taken into account. Some extracted parameters come with large uncertainties, which are due to the strong parameter correlations, though the underlying data are very precise. The predictions from the fitted LCDAs for $Q^{4}F_{1}(Q^{2})$ are confronted with those from the QCDSR parameters Braun:2006hz in Fig. 2(a). The former feature well-controlled uncertainties and a mild $Q^{2}$ dependence across the intermediate-to-large $Q^{2}$, in sharp contrast to the declining trends of the latter and in the previous QCDF calculations Huang:2024ugd ; Chen:2024fhj . It is obvious that the agreement between the theory and the data Sill:1992qw ; Christy:2021snt has been greatly improved.
@@
Figure 2: pQCD predictions for (a) the Dirac FF $Q^{4}F_{1}(Q^{2})$, (b) $Q^{4}F_{1}(Q^{2})$ from each twist combination $(t,t^{\prime})$, and (c) epep elastic scattering cross section, where “BLW 06” refers to the input of the QCDSR parameters Braun:2006hz .
@@
Figure 2(c) indicates that the pQCD prediction for the $ep$ cross section $\sigma$ matches perfectly the data in the range $5\leq Q^{2}\leq 30$ GeV2. The outcome from the QCDSR parameters fails to account for the observed $Q^{2}$ dependence apparently. Table 3 updates the pQCD predictions for the branching ratios (BRs), the direct CP asymmetries (CPAs), and the $S$- and $P$-wave CPAs in the hadronic heavy baryon decays $\Lambda_{b}\to p\pi$ and $pK$, whose theoretical uncertainties originate only from the variation of the $\Lambda_{b}$ baryon LCDA . The data and previous pQCD results Han:2024kgz based on the QCDSR parameters are quoted for comparison. We find that the predictions for both BRs and direct CPAs are highly sensitive to choices of the proton LCDAs, and that the updated pQCD calculations reveal better consistency with the $\Lambda_{b}\to pK^{-}$ data LHCb:2024iis .
Table 3: Updated pQCD predictions for the $\Lambda_{b}\to p\pi,pK$ decays.
$Br(10^{-6})$ $A_{CP}(\%)$ $A^{S}_{CP}(\%)$ $A^{P}_{CP}(\%)$
$\Lambda_{b}\to p\pi$
DataLHCb:2024iis $4.6\pm 0.8$ $0.2\pm 0.9$
pQCDHan:2024kgz $3.3^{+2.9}_{-1.7}$ $5^{+2}_{-3}$ $17^{+5}_{-9}$ $6^{+4}_{-5}$
this work $1.7^{+0.9}_{-0.7}$ $7^{+0}_{-7}$ $33^{+1}_{-3}$ $51^{+3}_{-3}$
$\Lambda_{b}\to pK$
DataLHCb:2024iis $5.5\pm 1.0$ $1.1\pm 0.8$
pQCDHan:2024kgz $2.8^{+3.3}_{-1.6}$ $6^{+3}_{-2}$ $5^{+5}_{-4}$ $21^{+39}_{-46}$
this work $4.4^{+2.6}_{-1.1}$ $3^{+2}_{-1}$ $7^{+1}_{-0}$ $66^{+7}_{-3}$

### v2

@@ Abstract
We present the first high-twist study of the proton form factors $F_{1,2}(Q^{2})$ in $ep$ elastic scattering based on the perturbative QCD $k_{T}$ factorization, $Q^{2}$ being momentum transfer squared. It is motivated by unexpectedly large higher-power contributions from subleading-twist light-cone distribution amplitudes (LCDAs), which are attributed to the enhancement in endpoint regions of parton momentum fractions. We highlight that the endpoint enhancement, tamed by the $k_{T}$ resummation effect, is crucial for accommodating the approximate scaling behavior of the $Q^{4}F_{1}(Q^{2})$ data at intermediate $Q^{2}\sim{\cal O}(10)$ GeV2. The proton LCDAs up to twist 6 are then extracted, and verified by the charge-parity asymmetries observed in hadronic heavy baryon decays. Our work provides new insights into the proton three-dimensional structure and manifests the precision requirement for reliable perturbative analyses of baryonic exclusive processes.
@@
We present the first pQCD study of the proton FFs at subleading powers, including the LCDAs up to twist 6 systematically. It will be shown that the endpoint enhancement maximizes the contributions to the FFs at combined twist 8, and the power expansion starts to converge above combined twist 9. The match of the pQCD predictions to the precise FF data from SLAC Arnold:1986nq ; Sill:1992qw ; Rock:1991jy and JLab Christy:2021snt selects the preferred parameters in the proton LCDAs, which confront the outcomes from QCD sum rules (QCDSR) Chernyak:1987nt , lattice QCD (LQCD) RQCD:2019hps and data-driven light-cone sum rules Anikin:2013aka . The determined proton LCDAs, which ought to be universal, are verified by the improved consistency between the predictions and the data for the charge-parity violation in the heavy baryon decay $\Lambda_{b}\to pK$ Han:2024kgz ; Han:2025tvc ; LHCb:2025ray . We highlight that the endpoint enhancement, tamed by the $k_{T}$ resummation effect, is crucial for explaining the scaling of $Q^{4}F_{1}(Q^{2})$, which cannot be achieved in the QCDF Huang:2024ugd ; Chen:2024fhj . Our work elaborates the necessity of the proton three-dimensional structure at high twists for understanding the $ep$ elastic scattering data.
@@
in which the plus (minus) component reads $P^{+}=E+Q/2$ ($P^{-}=E-Q/2$) with the proton energy $E=\sqrt{Q^{2}/4+m_{p}^{2}}$, and the parton momentum fractions $x_{i}^{(\prime)}$ and transverse momenta ${\bf k}^{(\prime)}_{iT}$ obey the conservation $\sum_{i}x^{(\prime)}_{i}=1$ and $\sum_{i}{\bf k}^{(\prime)}_{iT}=0$.
@@
The parton transverse momenta span three hierarchical scales: $Q$ (hard), $\sqrt{Q\Lambda_{\rm QCD}}$ (hard-collinear) and $\Lambda_{\rm QCD}$ (soft). The infrared logarithms associated with the soft scale are absorbed into the wave functions $\psi$. The large logarithms associated with the hard-collinear scale are resummed into the well-known Sudakov factors in the impact parameter space, and factored out of $\psi$, $\psi=e^{-S(x_{i},b_{i},Q)}\phi(x_{i},b_{i},\mu_{f})$, $\phi$ being the proton LCDAs. The explicit expression for the exponent $S$ have been given in Refs. Li:1992ce ; Botts:1989kf ; Li:1992nu . The scales $\mu_{r}=\mu_{f}$ for each diagram are set to the maximal virtuality of internal quarks and gluons. With the soft and hard-collinear dynamics being organized appropriately, Eq. (3) is applicable to exclusive processes at intermediate energies; the low $k_{T}$ (high $b$) region would be strongly suppressed by the Sudakov factors, such that internal particle virtualities are governed by finite transverse momenta in the endpoint region. We will ignore the intrinsic $b$ dependence Chai:2024tss and the factorization scale $\mu_{f}$ in $\phi(x_{i},b_{i},\mu_{f})$ for simplicity, so that the parton transverse distributions are completely described by the Sudakov factors.
@@
respectively, where $C$ stands for the charge conjugation, and $z$ is a light-like vector chosen in the minus direction. At NLO of the conformal spin expansion, five additional dimensionless parameters are related to local operators containing one derivative: $V_{1}^{d}$, $f_{1}^{d}$, $f_{2}^{d}$ ($A_{1}^{u}$, $f_{1}^{u}$) are associated with the derivatives on the $d$-quark ($u$-quark) fields in Eq. (4). The precise information on the above parameters are not yet available. Results from QCDSR Braun:2000kw ; Braun:2006hz and LQCD RQCD:2019hps , both acquired by the Regensburg group, exhibit magnificent discrepancies and uncertainties, even for the LO conformal spin parameters. Alternative methods for determining these parameters, such as data-driven analyses of the FFs, are worthwhile and can yield critical insights into the proton structure.
Table 1: Typical integrands of $F_{1}(Q^{2})$ from each combination of the initial (column) and final (row) proton LCDAs of various twists.
@@
Figure 1 presents the $Q^{2}$ dependencies of the various twist contributions to $Q^{4}F_{1}(Q^{2})$, as calculated using the QCDSR inputs for the nonperturbative parameters Braun:2006hz . For simplicity, only contributions from even twists are displayed. It turns out that the pieces from twist $n=8$ dominate, and the power expansion converges above $n=9$. Comparing Fig. 1 with the corresponding data in Fig. 2(a), we see that the predicted $Q^{2}$ dependence deviates dramatically from the observed scaling within $5\leq Q^{2}\leq 30$ GeV2, though the magnitudes of $Q^{4}F_{1}(Q^{2})$ are of the same order. We stress that the roughly constant $Q^{4}F_{1}(Q^{2})$ around intermediate $Q^{2}$ cannot be understood in collinear factorization, since the dominant $1/\tilde{Q}^{8}$ contribution leads to the descent of $Q^{4}F_{1}(Q^{2})$ with $Q^{2}$ definitely. Considering the additional hard-collinear scale $k_{T}$, which may modify the power-law behaviors associated with different twists, we examine this issue in the pQCD approach.
@@
If two parameters are always present as a product in the formula, the product is chosen as an element of $\left[a_{i}\right]$. The hard matrix ${\cal H}_{ij}$ depends only on $Q^{2}$, which has absorbed the scale evolution of the parameters. The pQCD predictions for the reduced $ep$ cross section at selected $Q_{l}^{2}$ points $\sigma_{l}^{\rm pQCD}\equiv\sigma^{\rm pQCD}(Q_{l}^{2})$ are obtained by inserting Eq. (5) into Christy:2021snt
@@
QCDSR $5.0\pm 0.5$ $2.7\pm 0.9$ $5.4\pm 1.9$ $0.23\pm 0.03$ $0.38\pm 0.15$ $0.4\pm 0.05$ $22\pm 0.05 0.0$$07\pm 0.05$
LQCD $3.67\pm 0.06$ $4.02\pm 0.38$ $8.37\pm 0.43$ $288\pm 0.007 0.0$$096\pm 0.010$
this work $3.67\pm 0.06$ $3.98\pm 0.37$ $8.32\pm 0.43$ $1.15\pm 0.44$ $1.27\pm 0.72$ $0.85\pm 0.12$ $0.59\pm 0.04$ $0.23\pm 0.03$
The parameters from the fit with $\chi^{2}/d.o.f.=1.6$ are compared with those from QCDSR Braun:2006hz and LQCD RQCD:2019hps in Table 2. The fit stabilizes the prior set of $f_{p}$, $\lambda_{1}$ and $\lambda_{2}$ derived in LQCD, including their uncertainties, confirming that these values are reasonable. However, significant shifts in $V_{1}^{d}$, $A_{1}^{u}$, $f_{1}^{u}$, $f_{1}^{d}$ and $f_{2}^{d}$ relative to QCDSR and LQCD results arppear. Because our pQCD study is conducted at LO in $\alpha_{s}$, the distinction can be conclusive after higher-order corrections are taken into account. Some extracted parameters come with large uncertainties, which are due to the strong parameter correlations, though the underlying data are very precise. The predictions from the fitted LCDAs for $Q^{4}F_{1}(Q^{2})$ are confronted with those from the QCDSR parameters Braun:2006hz in Fig. 2(a). The former feature well-controlled uncertainties and a mild $Q^{2}$ dependence across the intermediate-to-large $Q^{2}$, in sharp contrast to the declining trends of the latter and in the previous QCDF calculations Huang:2024ugd ; Chen:2024fhj . It is obvious that the agreement between the theory and the data Sill:1992qw ; Christy:2021snt has been greatly improved.
@@
Figure 2: pQCD predictions for (a) the Dirac FF $Q^{4}F_{1}(Q^{2})$, (b) $Q^{4}F_{1}(Q^{2})$ from each twist combination $(t,t^{\prime})$, and (c) epep elastic scattering cross section, where ”BLW 06” refers to the input of the QCDSR parameters Braun:2006hz .
@@
Figure 2(c) indicates that the pQCD prediction for the $ep$ cross section $\sigma$ matches perfectly the data in the range $5\leq Q^{2}\leq 30$ GeV2. The outcome from the QCDSR parameters fails to account for the observed $Q^{2}$ dependence apparently. At last, we verify the determined LCDAs by applying them to update the pQCD predictions for the $\Lambda_{b}\to pK$ decay. Based on their universality, we obtain a branching ratio of $Br=(4.4^{+2.6}_{-1.1})\times 10^{-6}$ and a direct CP asymmetry of $A_{CP}=(3^{+2}_{-1})\%$. The quoted uncertainties reflect only the dominant error source, which originates from the $\Lambda_{b}$ baryon LCDAs. Compared with the previous pQCD results, which relied on QCDSR parameters and yielded $Br=(2.8^{+3.3}_{-1.6})\times 10^{-6}$ and $A_{CP}=(-6^{+3}_{-2})\%$ Han:2024kgz , the updated calculations reveal better consistency with the experimental data $Br=(5.5\pm 1.0)\times 10^{-6}$ and $A_{CP}=(-1.1\pm 0.8)\%$ LHCb:2024iis .